OpenStudy (anonymous):

A local music store sells a popular brand of amplifier. In one month, they sold 56 units for a total profit of $2,917.60. If the shop sells 44 units the next month, how much profit will they make that month?

OpenStudy (anonymous):

x is the price per amp 56x = 2917.6 x = 2917.6/56 44 x will be 44x = (44*2917.6)/56

OpenStudy (anonymous):

Profit is proportional to the number of units sold so just use the ratio directly (44 / 56) * 2917.6 = answer